OpenStudy (anonymous):

d^2 r/dt^2 = (dx/dt)^2 *r solve this

OpenStudy (lgbasallote):

\[\huge \frac{d^2 r}{dt^2} = \left(\frac{dx}{dt}\right)^2 r\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

you sure? one is r and the other is x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@omjat u doin ODE or PDE?

OpenStudy (anonymous):

Ordinary Differential equations or Partial differential equations?

OpenStudy (anonymous):

@lgbasallote is right

OpenStudy (anonymous):

just give me answer